OpenStudy (anonymous):

Find the value of X: x^-2=1/16 4 and -4 only 4 16 and -16 only 16

OpenStudy (anonymous):

@alakara_kibo @Mathbreaker

OpenStudy (anonymous):

The answer is 4 and -4. Alakara is wrong, see you have (1/x)^2 = 1/16 now 1/x = sqrt(1/16) 1/x = +/- (1/4) now x = +/- 4 so it can be either 4 or -4